answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Suppose the variance is 64 Find the standard deviation?

Standard deviation is the square root of the variance; so if the variance is 64, the std dev is 8.

What is the probability of rolling a sum of 4?

For standard dice, there are 36 outcomes for rolling 2 dice. There are 3 ways to roll a 4; 3,1 1,3 & 2,2. The probability of sum of 4 is then 3/36 or 1/12 or 0.0833.

How many standard normal distributions are there?

Only one. A normal, or Gaussian distribution is completely defined by its mean and variance. The standard normal has mean = 0 and variance = 1. There is no other parameter, so no other source of variability.

If you throw a dice 6 times What is the chance you would get a 6 on?

First calculate the probability of NOT getting a six. This probability is 5/6 x 5/6 ... x 5/6, i.e., (5/6) to the power 6. Then subtract 1 minus this probability. This gives you the probability of getting at least one six.

How many different outcomes if you roll a dice?

6 Of course - assuming a cube die.

There are other dice available (for playing D&D) which can have for example 4 (tetrahedral dice) or 20 (icosahedral dice) outcomes.

If you roll a pair of six sided number cubes what is the probability of rolling two numbers whose sum is 6?

If you roll a pair of six sided number cubes the probability of rolling two numbers whose sum is 6 is 5/36 or 0.1389. There are 5 ways to get 6: 1,5 & 5,1 & 4,2 & 2,4 & 3,3. There are 36 outcomes for a roll of 2 dice.

If you draw one card from a normal deck of cards what is the probability of drawing a jack queen or king?

Probability of Jack being drawn is 4/52 since there are 4 Jacks and 52 cards in the deck. Also, the probability of drawing a Queen and King is 4/52. So, if you draw one card from a normal deck of cards the probability of drawing a jack or queen or king is 4/52 + 4/52 + 4/52 = 12/52 or 3/13 or 0.2308.

What is the probability of getting the sum of 5 when rolling 2 dice?

Probability of getting the sum of 5 when rolling 2 dice is 4/36 or 1/9 or 0.111. There are 4 ways to get a 5: 2,3 & 3,2 & 1,4 & 4,1. There are 36 outcomes on a roll of two dice.

In detailed way.

The possible elements in the sample space is got by 62 = 36 So n(S) = 36

THe event A is defined as getting sum of 5. The possible elements are (1,4) (2,3) (3,2) (4,1). Thus n(A) = 4

Required prob. = n(A)/n(S) = 4/36 = 1/9

If a die is rolled 3 times what's the chance of getting atleast one 6?

1/7776

50/50, or one chance in each set of three rolls. No matter how many times it's rolled, every roll is one in six. Just like flipping a coin. it's 50/50 heads or tails, even if you got 'heads' 20 times in a row the next flip will still be 50/50 chance of 'heads' again

What is the z score if 30 percent of the curve lies to the left of the z?

The z-score is 0.84. In the related link, look in the body of the table for .3 area (.2995 closest) and it yields the .84 z-score.

What is the meaning of Sampling distribution of the test statistic?

Given any sample size there are many samples of that size that can be drawn from the population. In the population is N and the sample size in n, then there are NCn, but remember that the population can be infinite.

A test statistic is a value that is calculated from only the observations in a sample (no unknown parameters are estimated). The value of the test statistic will change from sample to sample. The sampling distribution of a test statistic is the probability distribution function for all the values that the test statistic can take across all possible samples.

How many different ways can the letters of the word CHORD be arranged?

5 letters in first position there can be any one of the five, in the second position there are four possibilities left, in third position one of three remaining letters etc. 5 * 4 * 3 * 2 *1 = 120.

Range of probability?

If this is a homework assignment, please consider trying to answer it yourself first, otherwise the value of the reinforcement of the lesson offered by the assignment will be lost on you.

The range of probability for any event is between 0 and 1, inclusive.

If a coin is flipped and a die is rolled.what is the probability of getting heads and a 6?

The probability of flipping a heads is 1/2 and the probability of rolling a 6 is 1/6. By the laws of probability it would be logical to multiply them together, (1/2)(1/6) thus the answer being 1/12 with is roughly eight percent.